HIGH PRECISION IN AN IMPRECISE WORLD: THE IMPORTANCE OF THE MOLECULAR SPEC-TROSCOPY SYMPOSIUM TO THE ARMY

\underline{DAVID SKATRUD}, Army Research Office, Research Triangle Park, NC, 27709-2211.

The Army Research Office has been a consistent supporter of the OSU International Symposium on Molecular Spectroscopy for nearly 30 years. During this time, the Symposium has been of great value to the Army through a number of mechanisms. Not only have numerous specific research projects reported at the Symposium had major impacts on critical Army technologies; but in addition, an understanding of how to perform high-resolution spectroscopy research provides unique education and training on how to meaningfully model and explain many disparate complex phenomena. I will present historical examples, and also highlight related current Army research needs and thrusts.